OpenStudy (anonymous):

HELP !!!!!! Part 1: Create your own radical expression. Part 2: Convert it to a rational exponent. Part 3: Explain, in complete sentences, how the expression was converted.

OpenStudy (anonymous):

which part do you need help in ? :D

OpenStudy (anonymous):

all 3 lol

hero (hero):

Didn't I do this yesterday already?

hero (hero):

\[\sqrt[3]{27} = 27^{1/3} \] \[\sqrt[a]{b} = b^{1/a}\]

OpenStudy (anonymous):

I didnt put this up yesterday maybe you helped someone else